OpenStudy (anonymous):

A container has the following dimensions: 50 cm long, 30 cm wide, and 40 cm height; it has been filled with milk. a) How much milk in litters contains? b) Calculate the total weight of the milk. c) What is the pressure exerted over the bottom of the container if the density of the milk is 1 023kg/m3?

OpenStudy (anonymous):

\[1cm^3=1ml\] \[1/1000 liter = 1ml] density of milk: http://hypertextbook.com/facts/2002/AliciaNoelleJones.shtml w=mg = (density*volume)*g pressure = Force/Area = (mg)/A = (density*area*height*g)/area =density*height*g

OpenStudy (anonymous):

question a refers to the volume. Volume = length x width x height. So, V = (50)(30)(40). But this is in cubic centimeters, so you have to convert in in litters. (Google the conversion factor) Weight = mass*gravity. But you see, you need a value for mass (mass of the milk). To solve for the mass: mass = density*volume in this case, the density is given in question c. density = 1.023kg/m^3. Convert the volume in cubic meters, then, follow the equation. when you have the value of mass, (which should be in kg), solve for the weight. when solving for the pressure pressure = density*gravity*height where: density is 1.023kg/m^3 gravity is 9.8m/s^2 height/depth is 40 cm or .4m I hope this answers your question. Good luck!

OpenStudy (anonymous):

Thank you so much. So helpful your answer

OpenStudy (anonymous):

Your welcome!

OpenStudy (anonymous):

Can you help me with my other question? Im new on this.

OpenStudy (anonymous):

Actually, this is my first day here. Anyway, I'll check your other question and see if I can answer it.